Problem 7-5 (Part Level Submission)

Pina Company has four operating divisions. During the first quarter of 2017, the company reported aggregate income from operations of $212,300 and the following divisional results.
Division
I II III IV
Sales $254,000 $199,000 $501,000 $443,000
Cost of goods sold 204,000 190,000 301,000 247,000
Selling and administrative expenses 69,700 61,000 57,000 55,000
Income (loss) from operations $ (19,700) $ (52,000) $143,000 $141,000

Analysis reveals the following percentages of variable costs in each division.
I II III IV
Cost of goods sold 67 % 90 % 82 % 75 %
Selling and administrative expenses 39 62 50 58

Discontinuance of any division would save 50% of the fixed costs and expenses for that division.

Top management is very concerned about the unprofitable divisions (I and II). Consensus is that one or both of the divisions should be discontinued.

Collapse question part

(a)

Correct answer. Your answer is correct.
Compute the contribution margin for Divisions I and II. (Enter negative amounts using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).)
Division I Division II
Contribution margin $

Entry field with correct answer90137

$

Entry field with correct answer-9820

SHOW SOLUTION

SHOW ANSWER

LINK TO TEXT

Attempts: 2 of 15 used

Collapse question part

(b1)

Partially correct answer. Your answer is partially correct. Try again.
Prepare an incremental analysis concerning the possible discontinuance of Division I. (Round answers to 0 decimal places, e.g. 1525. If amount decreases net income then enter the amount using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).)
Continue Eliminate Net Income
Increase (Decrease)
Contribution margin $

Entry field with correct answer90137

$

Entry field with correct answer0

$

Entry field with correct answer-90137

Fixed costs
   Cost of goods sold

Entry field with correct answer67320

Entry field with correct answer33660

Entry field with correct answer33660

   Selling and administrative

42517Entry field with correct answer

Entry field with correct answer21258.5

Entry field with correct answer21258.5

      Total fixed expenses

Entry field with incorrect answer?

Entry field with correct answer54918.5

Entry field with correct answer54918.5

Income (loss) from operations $

?Entry field with incorrect answer

$

Entry field with incorrect answer?

$

-35218.50Entry field with correct answer

LINK TO TEXT
Answer #1
Continue Eliminate Net Income Increase (Decrease)
Contribution margin 90137 0 -90137
Fixed costs :
   Cost of goods sold 67320 33660 33660
   Selling and administrative 42517 21259 21259
      Total fixed expenses 109837 54919 54919
Income (loss) from operations -19700 -54919 -35219
Workings:
Continue Eliminate
Fixed costs :
   Cost of goods sold =(204000*33%) =67320*50%
   Selling and administrative =69700*61% =42517*50%